OpenStudy (anonymous):

Multiply and simplify: –4mp4 • (–4p) • (m3p) A. –16m4p6 B. –8m3p5 C. 16m3p5 D. 16m4p6

OpenStudy (anonymous):

D?

OpenStudy (mathstudent55):

\(–4mp^4 \times (–4p) \times (m^3p)\) Is this the problem?

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

Yes, D is correct.

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

\(–4mp^4 \times (–4p) \times (m^3p)\) Because of the commutative and associative properties, you can change the problem into this: \(= [-4 \times (-4)] \times (m \times m^3) \times (p^4 \times p \times p) \) \(=16m^4p^6\)

OpenStudy (mathstudent55):

You're welcome.